OpenStudy (helpingkid):

2/3 of the beads in a box are red. 1/4 are yellow and rest are blue. There are 42 more red beads blue beads. How many beads are there altogether?

OpenStudy (flcon04):

Total = 12x42/7 = 72

OpenStudy (cal.lavender):

r = number of red b = number of blue y = number of yellow r = b + 42 or b = r - 42 and I suppose r - b = 42 Fraction that are blue 1 - ( 2/3 + 1/4 ) 1 - ( 8/12 + 3/12 ) 1 - ( 11/12 ) b = 1/12 Since y is 1/4 and y = 3/12, y = 3b Since r is 2/3 and r = 8/12, r = 8b Combine r = b + 42 with r = 8b This means that b + 42 = 8b Subtract b from each side 42 = 7b 6 = b So we have b = 6 , y = 3b and r = 8b Since b + y + r = total By substituting 6 + 3(6) + 8(6) = total 6 + 18 + 48 = 72